OpenStudy (unklerhaukus):

Use transform definitions, and the evaluation of a suitable double integral, to calculate the Laplace transform of the following integral (ii)\[\mathcal L\left\{\int\limits_0^tf(t-u)g(u)\cdot\text du\right\}\]

OpenStudy (unklerhaukus):

\[\begin{align*}\mathcal L\left\{\int\limits_0^tf(t-u)g(u)\cdot\text du\right\}&=\int\limits_0^\infty\int\limits_0^tf(t-u)g(u)\cdot\text du\cdot e^{-pt}\text dt\\ &=\int\limits_0^\infty\int\limits_0^\infty f(t-u)g(u)H(t-u)\cdot\text du\cdot e^{-pt}\cdot\text dt\\ &=\int\limits_0^\infty\int\limits_0^\infty f(t-u)g(u)H(t-u)e^{-pt}\cdot\text dt\cdot\text du\\ &=\\ \end{align*} \]

OpenStudy (anonymous):

\(\begin{align*}\mathcal L\left\{\int\limits_0^tf(t-u)g(u)\cdot\text du\right\}&=\int\limits_0^\infty\int\limits_0^tf(t-u)g(u)\cdot\text du\cdot e^{-pt}\text dt\\ &=\int\limits_0^\infty e^{-pt}[\int \limits_0^tf(t-u)g(u)du]dt \\ &=\int\limits_0^\infty g(u)[\int \limits_u^\infty e^{-pt}f(t-u)dt]du \\ by \:changing \:the\:order\:of \:integration \end{align*}\) the region from u=0 to u=t is same as region from t=u to t=\(\infty\) now, substitute t-u = x in inner integral. then you will be able to split the integrals , one only containing t, and other only containing x....

OpenStudy (anonymous):

@UnkleRhaukus are u trying that ^ ?

OpenStudy (unklerhaukus):

yeah im working on it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

numbers ? which numbers ?

OpenStudy (unklerhaukus):

\[\begin{align*}\mathcal L\left\{\int\limits_0^tf(t-u)g(u)\cdot\text du\right\}&=\int\limits_0^\infty\int\limits_0^tf(t-u)g(u)\cdot\text du\cdot e^{-pt}\text dt\\ &=\int\limits_0^\infty\int\limits_0^\infty f(t-u)g(u)H(t-u)\cdot\text du\cdot e^{-pt}\cdot\text dt\\ &=\int\limits_0^\infty\int\limits_0^\infty f(t-u)g(u)H(t-u)e^{-pt}\cdot\text dt\cdot\text du\\ &=\int\limits_0^\infty g(u)\int\limits_0^\infty f(t-u)H(t-u)e^{-pt}\cdot\text dt\cdot\text du\\ \text{let }v=t-u&\\ \text dv =\text dt\\ t=0\rightarrow v=-u\\ t=\infty \rightarrow v=\infty\\ &=\int\limits_0^\infty g(u)\int\limits_{-u}^\infty f(v)H(v)e^{-p(v+u)}\cdot\text dv\cdot\text du\\ &=\int\limits_0^\infty g(u)\int\limits_{-u}^\infty f(v)e^{-p(v+u)}\cdot\text dv\cdot\text du\\ \end{align*}\]

OpenStudy (anonymous):

didn't change the order of integration.....? plus you H is throwing me of...why(and how) its used there ... ?

OpenStudy (unklerhaukus):

im not sure wether or not the unsure the unit heaviside step function

OpenStudy (anonymous):

if u follow (and understand) what i am suggesting, u don't need H

OpenStudy (unklerhaukus):

\[\begin{align*}\mathcal L\left\{\int\limits_0^tf(t-u)g(u)\cdot\text du\right\} &=\int\limits_0^\infty\int\limits_0^t f(t-u)g(u)\cdot\text du\cdot e^{-pt}\cdot\text dt\\ &=\int\limits_0^\infty e^{-pt}\int\limits_0^t f(t-u)g(u)\cdot\text du\cdot\text dt\\ &=\int\limits_0^\infty e^{-pt}\int\limits_0^t f(t-u)g(u)\cdot\text dt\cdot\text du\\ \text{let }v=t-u&\\ \text dv =\text dt\\ t=0\rightarrow v=-u\\ t=\infty \rightarrow v=\infty\\ &=\int\limits_0^\infty g(u)\int\limits_{u}^\infty e^{-p(v+u)}f(v)\cdot\text dv\cdot\text du\\ &=\\ \end{align*}\]

OpenStudy (anonymous):

still, didn't change the order of integration.......

OpenStudy (anonymous):

u=0 to u=t cahnges to t=u to t=infinity

OpenStudy (anonymous):

then put t-u=v

OpenStudy (unklerhaukus):

this question is confusing me

OpenStudy (anonymous):

whats the confusion about ? changing order of integration ?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):